### v1

@@ Abstract
Quantifying sarcomere structure organization in human-induced pluripotent stem cell-derived cardiomyocytes (hiPSC-CMs) is crucial for understanding cardiac disease pathology, improving drug screening, and advancing regenerative medicine. Traditional methods, such as manual annotation and Fourier transform analysis, are labor-intensive, error-prone, and lack high-throughput capabilities. In this study, we present a novel deep learning-based framework that leverages cell images and integrates cell features to automatically evaluate the sarcomere structure of hiPSC-CMs from the onset of differentiation. This framework overcomes the limitations of traditional methods through automated, high-throughput analysis, providing consistent, reliable results while accurately detecting complex sarcomere patterns across diverse samples. The proposed framework contains the SarcNet, a linear layers-added ResNet-18 module, to output a continuous score ranging from one to five that captures the level of sarcomere structure organization. It is trained and validated on an open-source dataset of hiPSC-CMs images with the endogenously GFP-tagged alpha-actinin-2 structure developed by the Allen Institute for Cell Science (AICS). SarcNet achieves a Spearman correlation of 0.831 with expert evaluations, demonstrating superior performance and an improvement of 0.075 over the current state-of-the-art approach, which uses linear regression. Our results also show a consistent pattern of increasing organization from day 18 to day 32 of differentiation, aligning with expert evaluations. By integrating the quantitative features calculated directly from the images with the visual features learned during the deep learning model, our framework offers a more comprehensive and accurate assessment, enhancing the further utility of hiPSC-CMs in medical research and therapy development.
keywords:
hiPSC-CMs, sarcomere structure organization, SarcNet, linear regression, deep learning.
@@ 1 Introduction
Human-induced pluripotent stem cell-derived cardiomyocytes (hiPSC-CMs) are cells safely collected from the blood or skin of living humans and then reprogrammed into human heart muscle cells [9]. Fully characterized hiPSC-CMs have a range of critical applications, including disease modeling, drug discovery, regenerative and precision medicine [3, 10, 12]. Specifically, these cells have attracted considerable attention as a promising alternative for modeling arrhythmogenic disorders and assessing the cardiac function of patients [24]. In another instance, Vicente et al. (2018) introduced the Comprehensive in vitro proarrhythmia Assay (CiPA) as a method to assess the impact of drugs on various ion channels in hiPSC-CMs and to predict the risk of proarrhythmia [22].
To effectively utilize these applications, hiPSC-CMs must acquire key features and capacities that accurately reflect the electrical activity, calcium dynamics, structure, and contractility of mature cardiomyocytes [3, 4]. During the differentiation of hiPSC-CMs, gene expression varies, with immature cardiomyocytes showing higher levels of the TNNI1 gene, while mature cells primarily express TNNI3 [2, 8]. Therefore, it is necessary to establish a framework to quantify hiPSC-CM maturation.
@@
Quantitative approaches for tracking hiPSC-CMs development are vital but currently restricted [3]. Traditional methods for assessing cardiomyocyte structures include manual annotation, which is labor-intensive and lacks high-throughput capabilities, and Fourier transform analysis on hiPSC-CMs alignment, which is error-prone on images with thicker multi-cell layer [15]. Recently, various methodologies have been employed to assess the maturation of hiPSC-CMs, focusing on distinct characteristics, such as sarcomere configuration, electrophysiological attributes, metabolism, and gene expression profiles [1]. For example, measuring the TNNI3 to TNNI1 ratio is one way to quantify hiPSC-CMs based on gene expression [2]. The structure of the sarcomere can be evaluated using hiPSC-CM imaging, an important technique to consider. This is particularly relevant because the contraction capabilities of adult cardiomyocytes are closely tied to the organization and structure of the sarcomeres and myofilaments [21]. In particular, mature hiPSC-CMs are notably longer and display a higher degree of structural organization compared to their immature counterparts [17].
@@
Several studies have employed artificial intelligence (AI) to quantify the maturity of hiPSC-CMs by analyzing fluorescent images [19, 7]. The authors in [19] (2015) first defined a set of 11 metrics to capture the increasing organization of sarcomeres within striated muscle cells during their developmental process and then used machine learning algorithms to score the phenotypic maturity of hiPSC-CMs unbiasedly. However, due to the limited availability of hiPSC-CMs at various developmental stages, the model was trained on primary cardiomyocytes from neonate rats (rpCMs) and was tested on immature hiPSC-CMs only. In [7] published in 2021, Gerbin, Kaytlyn A., et al. used linear regression to distinguish stages of myofibrillar organization in hiPSC-CMs at the single-cell level. They utilized 11 cell features, six of which originated from deep learning. The model reached a Spearman correlation of 0.63 and 0.67 on two test sets. Despite these advances, translation to clinical application remains a distant goal.
@@
To address the gap, in this study, we propose SarcNet, a cell-features concatenated and linear layers-added ResNet-18 convolutional neural network (CNN) for automatically quantifying sarcomere structure organizations on single-cell images of hiPSC-CMs. In particular, this model leverages predictions by concatenating the output from the ResNet-18 module and a representation vector of quantitative single-cell measurements of subcellular organization, followed by the addition of four linear layers to gain deeper linear representations. To evaluate the proposed framework, we compare SarcNet with a combined model, which integrates predictions from linear regression and the ResNet-18 module to generate the final score. SarcNet obtains a higher performance, with a Spearman correlation of 0.831. Furthermore, we assess the performance of each approach using two different feature extraction protocols as described in Section 2.3.1. While the performance of the combined model drops significantly with Protocol 1, our SarcNet model maintains nearly consistent performance. This result indicates that using SarcNet, the features extracted by Protocol 1 can be sufficient for scoring sarcomere structure organizations without a complicated deep learning-based feature extraction method as in Protocol 2. In addition, SarcNet also outperforms linear regression and other state-of-the-art CNN models. Finally, we evaluate the proposed model using Gradient-weighted Class Activation Mapping (Grad-CAM) [20], an explanatory algorithm. The heatmaps generated by Grad-CAM reveal that the important regions in the correctly predicted images correspond to the majority of sarcomere patterns within the cell. Our main contributions through this work can be summarized as follows:
@@
    We introduce a novel AI-based framework to effectively analyze and score sarcomere organizations in fluorescently tagged hiPSC-CMs single cell images by integrating feature extraction with the existing state-of-the-art CNN model.
@@
    We conduct thorough experiments to evaluate the efficacy of the proposed solution and compare it to other methods. The experimental findings demonstrated that our SarcNet model with 11 features extracted by Protocol 2 enhances performance by at least 0.075 and 0.022 for the Spearman correlation metric compared to linear regression and the combined model. Notably, using only five features from Protocol 1 does not significantly affect the model performance.
@@
    We can speed up the training and inference by simplifying the feature extraction process and decreasing the number of features while maintaining performance. Our codes and models will be released at https://github.com/vinuni-vishc/sarcnet.
The remainder of the paper is organized as follows. Section 2 describes an overview of the proposed framework for quantifying sarcomere structure organization. In Section 3, we assess the databases and discuss implementation specifics. Section 4 presents the experiment results, visualizations, and limitations. Section 5 concludes the paper.
@@ 2 Methodology
This section introduces details of the proposed approach. We first give an overview of our framework for predicting a continuous score of sarcomere structure organization on single-cell imaging of hiPSC-CMs (Section 2.1). We then provide a formulation of the problem as a regression task (Section 2.2). Next, the framework architecture is described (Section 2.3).
@@ 2.1 Overall Framework
Figure 1: An illustration of our overall framework, which aims to quantify sarcomere structure organization. The system takes the hiPSC-CM images as input and outputs the scores of alpha-actinin-2 patterns, ranging from one to five. For feature extraction, three sets with a total of 11 cell features can be extracted using the techniques described in [7], including CellProfiler-based feature extraction, Gray-Level Co-occurrence Matrix-based (GLCM-based) feature extraction and deep learning-based feature extraction. Our model also localizes the most relevant portions of the image using a heatmap generated by the Grad-CAM approach and proclaims the increasing organization between day 18 and day 32 using the histogram.
Fig. 1 provides an overview of the proposed approach, which is a regression framework using CNN. It predicts a continuous score of sarcomere structure organization by taking fluorescent images of hiPSC-CMs single cells as input. To train the proposed model, an expert-scored hiPSC-CMs single cells dataset of 5,761 images has been used (Section 3.1). We enhance the prediction performance by integrating feature extraction into the model architecture. Model performance is evaluated using four metrics: Spearman correlation, mean absolute error (MAE), mean squared error (MSE), and $R^{2}$ score (Section 3.3). A visual explanation module based on Grad-CAMs is also used for model interpretation. Last but not least, we confirm the pattern of increasing maturity levels of sarcomere structure between day 18 and day 32-time points.
2.2 Problem Formulation
In a regression problem setting, we are given a training set $\mathcal{D}$ consisting of N samples, $\mathcal{D}$ = {$(x^{(i)},y^{(i)}$); i = 1,…, N} where each input image $x^{(i)}\in\mathcal{X}$ is associated with a continuous value $y^{(i)}\in\mathcal{Y}$. Our goal is to approximate a mapping function $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$ from input images, which are fluorescently tagged single-cell images of hiPSC-CMs to continuous output variables representing the organization of sarcomere structures. In general, this learning task could be performed by training a CNN, parameterized by weight $\theta$ that the MSE loss function is minimized over the training set $\mathcal{D}$. As the desired output is a continuous quantity, the activation function at the output layer is linear. The loss function MSE is given by
$\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}|y^{(i)}-\hat{y}^{(i)}|^{2},$ (1)
where $y^{(i)}$ is the ground truth and $\hat{y}^{(i)}$ is the value predicted by the network.
@@ 2.3.1 Feature Extraction
Three sets with a total of 11 cell features can be extracted for each input image using techniques provided in [7]. More details, Set 1, the CellProfiler-based feature extraction [18], generates two cell morphology measurement metrics: cell area and aspect ratio. Set 2, the GLCM-based method, produces three metrics describing sarcomere alignment: max coefficient of variation, peak height, and peak distance. Finally, Set 3, the deep learning-based method, provides six additional metrics representing the fraction of the cell area covered by each class: background, diffuse/other, fibers, disorganized puncta, organized puncta, and organized z-discs. The features extracted from Set 1 and Set 2 together are referred to as Protocol 1. The features extracted from Set 1, Set 2, and Set 3 together are referred to as Protocol 2.
@@ 2.3.2 ResNet-18 Module
ResNet-18 [11], one of the state-of-the-art image classification models, is adapted for the single-cell image classification task. It receives input images of size $3\times 224\times 224$ and generates labels ranging from one to five, corresponding to the score of sarcomere structure organization. The ResNet-18 module architecture used in this study consists of 18 convolution layers, an average pooling layer, and three linear layers. Compared to the original architecture, two fully connected layers are added at the end before outputting the result.
@@ 2.3.3 SarcNet
Figure 2: SarcNet architecture for scoring sarcomere structure organization, which are continuous values ranging from one to five.
In this study, we propose the SarcNet, which comprises two branches of the ResNet-18 module and linear layers as depicted in Fig. 2. First, the former requires an input image with a size of $3\times 224\times 224$ to pass through the ResNet-18 module (described in 2.3 - ResNet-18 module). Regarding the second branch, the input feature vector would be passed through three linear layers with rectified linear unit (ReLU) activation functions to extract more deep linear representations. After concatenating with the output from the ResNet-18 module, a new representation vector is created, consisting of the information contributed equally from both input images and input features. This vector is then processed through four linear layers to get the final score.
@@ 3.1.1 Dataset
In this study, we use a publicly available dataset of alpha-actinin-2-mEGFP-tagged hPSC-CMs single cells at days 18 and 32 time points since the initiation of differentiation, provided by the Allen Institute for Cell Science (AICS) [7]. The dataset consists of two components: images of single cells and tabular data containing cell features, along with the overall alpha-actinin-2 organization score corresponding to each image.
@@
As described [7], each cell is manually scored for the structural maturity of its sarcomere organization by two experts; thus, all cells are assessed twice. Based on the majority of their organization, experts categorize cells into five score groups, from one to five. A score of 1 indicates cells with sparse, disorganized puncta; a score of 2 indicates cells with denser, more organized puncta; a score of 3 indicates cells with a combination of puncta and other structures like fibers and z-discs; a score of 4 indicates cells with regular but not aligned z-discs and finally a score of 5 indicates cells with almost organized and aligned z-discs. Cells lacking alpha-actinin-2 mEGFP protein expression were assigned a score of “0” and subsequently excluded from further analysis. After filtering, there are a total of 5,761 images of different sizes. The number of cell images in each score category is shown in Table 1. For consistency, we define ground truth (GT) in our downstream analysis as the mean score from two experts.
Table 1: Number of cells in each score category
Score 1 Score 2 Score 3 Score 4 Score 5
Expert 1 293 708 3,868 798 94
Expert 2 115 1,007 2,979 1,527 113
@@
The tabular data component provides cell features extracted by CellProfiler, GLCM, and deep-learning model. In this work, we utilize a subset of 11 features from the tabular data in the dataset package, including cell area, cell aspect ratio, max coefficient of variation, peak height, peak distance, fraction cell area background, fraction cell area diffuse/other, fraction cell area fibers, fraction cell area disorganized puncta, fraction cell area organized puncta, and fraction cell area organized z-discs. Table 2 shows a statistical summary of the mean and standard deviation (SD) of these 11 metrics. It is evident that cell area, max coefficient of variation, and peak distance increase as cells mature. Also, the fraction cell area background is highest in cells scored as 1, while the fraction cell area organized z-disc is highest in cells scored as 5. These features are utilized differently in the two protocols: Protocol 1 employs the first five features, whereas Protocol 2 incorporates all 11 features, as depicted in Section 2.3.1.
Table 2: Statistical summary of cell features in each score category
Mean (SD) 1.0 1.5 2.0 2.5 3.0 3.5 4.0 4.5 5.0 Count 81 234 440 622 2,543 1,108 572 120 41 Cell area 1,732 (1,040) 1,620 (1,016) 1,925 (1,070) 1,923 (993) 2,272 (1,207) 2,557 (1,489) 3,090 (2,086) 2,669 (1,394) 2,717 (1,300) Set 1 Cell aspect ratio 0.49 (0.16) 0.50 (0.18) 0.48 (0.19) 0.50 (0.19) 0.50 (0.18) 0.54 (0.18) 0.50 (0.18) 0.46 (0.18) 0.46 (0.19) Max coefficient of variation 0.26 (0.16) 0.25 (0.16) 0.24 (0.14) 0.25 (0.14) 0.30 (0.15) 0.32 (0.15) 0.37 (0.16) 0.49 (0.16) 0.57 (0.17) Peak height 0.47 (0.09) 0.48 (0.10) 0.44 (0.09) 0.44 (0.09) 0.44 (0.10) 0.44 (0.10) 0.46 (0.10) 0.55 (0.08) 0.60 (0.07) Set 2 Peak distance 1.50 (0.50) 1.72 (0.41) 1.69 (0.31) 1.72 (0.25) 1.76 (0.19) 1.77 (0.12) 1.78 (0.10) 1.79 (0.08) 1.79 (0.06) Fraction cell area background 0.20 (0.28) 0.10 (0.19) 0.04 (0.11) 0.04 (0.11) 0.03 (0.09) 0.02 (0.08) 0.02 (0.06) 0.04 (0.08) 0.05 (0.10) Fraction cell area diffuse/other 0.04 (0.06) 0.09 (0.09) 0.14 (0.11) 0.15 (0.12) 0.19 (0.14) 0.18 (0.14) 0.11 (0.10) 0.11 (0.09) 0.08 (0.07) Fraction cell area fiber 0.30 (0.22) 0.34 (0.20) 0.34 (0.19) 0.32 (0.19) 0.27 (0.18) 0.14 (0.12) 0.09 (0.10) 0.08 (0.09) 0.06 (0.07) Fraction cell area disorganized puncta 0.09 (0.11) 0.13 (0.13) 0.20 (0.15) 0.15 (0.13) 0.08 (0.09) 0.03 (0.06) 0.01 (0.02) 0.01 (0.02) 0.01 (0.02) Fraction cell area organized puncta 0.35 (0.23) 0.34 (0.20) 0.27 (0.17) 0.30 (0.20) 0.30 (0.18) 0.29 (0.18) 0.26 (0.16) 0.20 (0.13) 0.16 (0.09) Set 3 Fraction cell area organized z-discs 0.01 (0.02) 0.01 (0.02) 0.01 (0.03) 0.03 (0.05) 0.13 (0.14) 0.34 (0.19) 0.49 (0.20) 0.56 (0.19) 0.63 (0.19)
3.1.2 Settings

### v2

@@ Abstract
Quantifying sarcomere structure organization in human-induced pluripotent stem cell-derived cardiomyocytes (hiPSC-CMs) is crucial for understanding cardiac disease pathology, improving drug screening, and advancing regenerative medicine. Traditional methods, such as manual annotation and Fourier transform analysis, are labor-intensive, error-prone, and lack high-throughput capabilities. This paper presents a novel deep learning-based framework that leverages cell images and integrates cell features to automatically evaluate the sarcomere structure of hiPSC-CMs from the onset of differentiation. The proposed framework contains the SarcNet, a cell-features concatenated and linear layers-added ResNet-18 module, to output a continuous score ranging from 1 to 5 that captures the level of sarcomere structural organization. SarcNet achieves a Spearman correlation of 0.831 with expert evaluations, demonstrating superior performance and an improvement of 7.5% over the previous approach, which uses linear regression. Our results also show a consistent pattern of increasing organization from day 18 to day 32 of differentiation, aligning with expert evaluations. This study contributes to the development of a new automated tool for quantifying sarcomere structural organization in hiPSC-CMs, advancing cardiac research.
Index Terms—  hiPSC-CMs, sarcomere, deep learning.
@@ 1 Introduction
Human-induced pluripotent stem cell-derived cardiomyocytes (hiPSC-CMs) are cells safely collected from the blood or skin of living humans and then reprogrammed into human heart muscle cells [1]. Fully characterized hiPSC-CMs have a range of critical applications, including disease modeling, drug discovery, and precision medicine [2]. Specifically, these cells have attracted considerable attention as a promising alternative for modeling arrhythmogenic disorders and assessing the cardiac function of patients [3]. In another instance, Vicente et al. [4] introduced the Comprehensive in vitro proarrhythmia Assay (CiPA) as a method to assess the impact of drugs on various ion channels in hiPSC-CMs and to predict the risk of proarrhythmia.
@@
To effectively utilize these applications, hiPSC-CMs must accurately reflect the electrical activity, calcium dynamics, structure, and contractility of mature cardiomyocytes [2]. Recently, various methods have been employed to assess the maturation of hiPSC-CMs, using a diversity of characteristics, such as sarcomere configuration, electrophysiological attributes, metabolism, and gene expression profiles [5, 6]. Among them, evaluating the structure of the sarcomere, the basic contractile unit of hiPSC-CMs, based on imaging data is an important technique to consider. The reason is that the organization and structure of the sarcomeres and myofilaments are closely tied to the contraction capabilities of adult cardiomyocytes [7]. In particular, mature hiPSC-CMs are notably longer and display a higher degree of structural organization compared to their immature counterparts [8].
@@
Several studies have employed artificial intelligence (AI) to quantify maturity of sarcomere structure in hiPSC-CMs by analyzing fluorescent images [9, 10]. Pasqualini et al. [9] first defined a set of 11 metrics to capture the increasing organization of sarcomeres within striated muscle cells during their developmental process and then used machine learning algorithms to score the phenotypic maturity of hiPSC-CMs unbiasedly. However, due to the limited availability of hiPSC-CMs at various developmental stages, the model was trained on primary cardiomyocytes from neonate rats (rpCMs) and was tested on immature hiPSC-CMs only. Gerbin et al. [10] used linear regression with 11 cell features to distinguish stages of myofibrillar organization in hiPSC-CMs. The model reached a Spearman correlation of 0.63 and 0.67 on two testing sets, respectively. Despite these advances, further development is needed since clinical applications require higher performance.
@@
To address the gap, we propose in this study SarcNet, a cell-features concatenated and linear layers-added ResNet-18 convolutional neural network (CNN) [11] for automatically quantifying sarcomere structural organizations on single-cell images of hiPSC-CMs. In particular, the proposed model leverages predictions by concatenating the output from the ResNet-18 module and a representation vector of quantitative single-cell measurements of subcellular organization. In addition, we assess the performance of the proposed framework using two different feature extraction protocols. We further assess the explainability of the proposed model by utilizing Gradient-weighted Class Activation Mapping (Grad-CAM) [12]. Our main contributions can be summarized as follows:
@@
    We introduce a novel AI-based framework to effectively score sarcomere organizations in fluorescently tagged hiPSC-CMs single cell images by integrating feature extraction with the state-of-the-art CNN model.
@@
    We conduct thorough experiments to evaluate the efficacy of the proposed solution. The experimental findings demonstrated that SarcNet model enhances performance by 7.5% for Spearman correlation metric compared to linear regression, the previously used method. Notably, we can speed up the training and inference by simplifying the feature extraction process while maintaining performance.
@@
    Our codes are released at https://github.com/vinuni-vishc/sarcnet.
The rest of the paper covers the proposed approach (Section 2), the experiments and results (Section 3), and concludes with Section 4.
@@ 2 METHODOLOGY
2.1 Problem Formulation
In a regression problem setting, we are given a training set $\mathcal{D}$ consisting of N samples, $\mathcal{D}$ = {$(x^{(i)},y^{(i)}$); (i) = 1,…, N} where each input image $x^{(i)}\in\mathcal{X}$ is associated with a continuous value $y^{(i)}\in\mathcal{Y}$. Our goal is to approximate a mapping function $f_{(\theta)}:\mathcal{X}\rightarrow\mathcal{Y}$ from input images, which are fluorescently tagged single-cell images of hiPSC-CMs to continuous output variables representing its sarcomere structural organization. This learning task could be performed by training a CNN that minimize the MSE loss function, which is given by
$\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}|y^{(i)}-\hat{y}^{(i)}|^{2},$ (1)
where $y^{(i)}$ is the ground truth and $\hat{y}^{(i)}$ is the predicted value.
@@ 2.2 Overall Framework
Fig. 1: An illustration of the proposed overall framework, which aims to quantify sarcomere structure organization. The system takes the hiPSC-CM images as input and outputs the scores of alpha-actinin-2 patterns, ranging from 1 to 5. Grad-CAM heatmaps is used to highlight important visual features.
Fig. 1 provides an overview of the proposed approach, which is a regression framework using CNN. It predicts a continuous score of sarcomere structure organization by taking fluorescent images of hiPSC-CMs single cells as input. We enhance the prediction performance by integrating feature extraction into the training procedure. Model performance is evaluated using four metrics: Spearman correlation, mean absolute error (MAE), mean squared error (MSE), and $R^{2}$ score. A visual explanation module based on Grad-CAMs is also used for model interpretation. We also examine the changes in cell organization over time and confirm the increasing maturity of sarcomere structures between day 18 and day 32 time points.
@@ 2.3.1 Feature Extraction
Three sets with a total of 11 cell features are extracted for each input image using techniques provided in [10]. More details, Set 1, the CellProfiler-based [13] feature extraction generates two cell morphology measurement metrics: cell area and aspect ratio. Set 2, the GLCM-based [14] method produces three metrics describing sarcomere alignment: max coefficient of variation, peak height, and peak distance. Finally, Set 3, the deep learning-based method provides six additional metrics representing the fraction of the cell area covered by each class: background, diffuse/other, fibers, disorganized puncta, organized puncta, and organized z-discs. The features extracted from Set 1 and Set 2 together are referred to as Protocol 1. The features extracted from Set 1, Set 2, and Set 3 together are referred to as Protocol 2.
@@ 2.3.2 ResNet-18 Module
ResNet-18 [11], one of the state-of-the-art image classification models, is adapted for the sarcomere structural level prediction task as the deep residual learning framework makes it well-suited for capturing the intricate hierarchical patterns of sarcomere structures. The ResNet-18 module architecture used in this study consists of 18 convolution layers, an average pooling layer, and three linear layers. Compared to the original architecture, two fully connected layers are added at the end before outputting the result.
@@ 2.3.3 SarcNet
The proposed SarcNet framework comprises two branches of the ResNet-18 module and linear layers as depicted in Fig. 1. First, the former requires an input image with a size of $3\times 224\times 224$ to pass through the ResNet-18 module (described in 2.3). Second, the input feature vector would be passed through three linear layers with rectified linear unit (ReLU) activation functions to extract more deep linear representations. After concatenating with the output from the ResNet-18 module, a new representation vector is created, consisting of the information contributed equally from both input images and input features. This vector is then processed through four linear layers to get the final score.
@@ 3.1.1 Dataset
In this study, we use a publicly available dataset of alpha-actinin-2-mEGFP-tagged hPSC-CMs single cells at days 18 and 32 time points since the initiation of differentiation, provided by the Allen Institute for Cell Science (AICS) [10]. The dataset consists of two components: images of single cells and tabular data containing cell features, along with the alpha-actinin-2 organization score corresponding to each image.
@@
As described in the original paper [10], each cell is manually scored for the structural maturity of its sarcomere organization by two experts. Based on the majority of their organization, experts categorize cells into five score groups, from 1 to 5. A score of 1 indicates cells with sparse, disorganized puncta; a score of 2 indicates cells with denser, more organized puncta; a score of 3 indicates cells with a combination of puncta and other structures like fibers and z-discs; a score of 4 indicates cells with regular but not aligned z-discs and finally a score of 5 indicates cells with almost organized and aligned z-discs. Cells lacking alpha-actinin-2 mEGFP protein expression were assigned a score of “0” and subsequently excluded from further analysis. After filtering, there are a total of 5,761 images of different sizes. Expert 1 categorized 293 images as Score 1, 708 as Score 2, 3,868 as Score 3, 798 as Score 4, and 94 as Score 5. Expert 2 classified 115 images as Score 1, 1,007 as Score 2, 2,979 as Score 3, 1,527 as Score 4, and 113 as Score 5. For consistency, we define ground truth in downstream analysis as the average score from two experts.
@@
The tabular data component provides cell features extracted by CellProfiler, GLCM, and deep-learning model. 11 features from the tabular data in the dataset package are utilized, including cell area, cell aspect ratio, max coefficient of variation, peak height, peak distance, fraction of cell area for background, diffuse/other, fibers, disorganized puncta, organized puncta, and organized z-discs. These features are applied into two protocols as depicted in Section 2.3.1.
3.1.2 Experimental Settings